OpenStudy (anonymous):

x+y=6 where do the lines intersect?

OpenStudy (anonymous):

at the origin which is 0

OpenStudy (anonymous):

oh. How did you find that?

OpenStudy (anonymous):

set x=0 to solve for y and set y=0 to solve for x,

OpenStudy (anonymous):

then u can get the points of (x,y)

OpenStudy (anonymous):

so like on x+2y=0... Would I do like 0+2y=0 making x -2 and like x+0y=0? (-2,0)?

OpenStudy (anonymous):

set y=0,to solve for x x+0=6 x=6 when x=6, 6+y=6 y=6-6 y=0

OpenStudy (anonymous):

okay sooo like with x + y = -3 x+0=-3 x=-3 when x = -3 -3+y=-3 y=-3-3 y=-6? maybe?

OpenStudy (anonymous):

yeah,u'r right

OpenStudy (anonymous):

yaayyy thank you sooo much! :D

OpenStudy (anonymous):

welcome :-)